OpenStudy (anonymous):

Please Help!!! solve by elimination 3x+2y=1 4x+3y=-2

OpenStudy (anonymous):

(3x+2y=1)3 (4x+3y=-2)-2 9x+6y=3 -8x-6y=4 Add them together, and the y's cancel. Solve for y, then solve for x